OpenStudy (anonymous):

all multiples of odd numbers are odd, is this statement true? justify your answer

OpenStudy (math&ing001):

Hello there, not all multiples of odd numbers are odd, as a proof you can multiply any odd number by 2 and you'll get an even number. Example : 7*2=14

OpenStudy (anonymous):

This was answered don't bump it

OpenStudy (anonymous):

If you need a proof let any odd be 2k+1 and 2j+1 next multiple (2k+1)(2j+1) = 4kj+2k+2j+1 = 2(kj+k+j) +1 = 2m+1 which is odd So assuming you mean what do you get when you multiply two odd numbers then yes it is odd. Otherwise math&ing001 is correct.